OpenStudy (anonymous):

Homework 11, Part1 question c) For f=740k L=0.65mH RL=4ohm RC=490k It's the only question left, I tried values from 4 to 10, but it wont check, can't get a working formula either... Any help? :)

OpenStudy (anonymous):

Hello C = 71.1645106214 pF Band = 5.54357213286 kHz If you second part has the value f=1180kHz C = 27.98 pF Band = 12.58kHz If you got green marks, please click in "best response" :D

OpenStudy (anonymous):

hi plz tell me the formula for bandwidth in h11p1

OpenStudy (anonymous):

For the bandwidth: dw = RL/L + 1/(Rc *C) after that, find f=(w/2*PI)/(1000) If works, please click at "Best response" :D

OpenStudy (anonymous):

hey it's not working my values are RL≈4.0Ω, L≈0.65mH, RC≈490.0kΩ, c=46pF, f=920.0kHz & for another one 3 bit f=1030.0kHz & c=36pF plzz. give me answer...

OpenStudy (anonymous):

Hello gurukandula I am not sure if this values are correct. Please, check if its ok :) For 920KHz C = 46.04 pF Bandwidth = 8.03428050596 KHz For 1030KHz C = 36.73 pF Bandwidth = 9.82248626529 KHz If works, please click at "Best Response"